# Gradient Interference-Aware Graph Coloring for Multitask Learning

### v1

@@ Appendix G Computational Complexity of One Refresh (and Amortized Over Training)
We analyze the computational and memory complexity of the proposed interference-aware scheduler per refresh and its amortized cost over training. The former accounts for the cost of a single refresh operation while the latter represents the average cost distributed across all training steps. We distinguish the work required by the underlying mutli-task training objective (.e.g, backpropagatoin to obtain gradients) from the scheduler overhead (EMA maintenance, cosine computation, conflict graph construction, and color).
G.1 Notation
@@
G.2 Per-refresh complexity (time and space)
@@
$m_{i}\leftarrow\beta m_{i}+(1-\beta)g_{i}$ (G.2)
@@
$\widehat{C}=\widetilde{M}\widetilde{M}^{\top}$ (G.3)
where $\widetilde{M}$ is the row-normalized version of $M$. It thresholds $\widehat{C}$ at $\tau$ to obtain the conflict adjacency. Finally, it applies a graph-coloring routine to the resulting simple graph ([40]).
@@
The construction of $\widehat{C}$ proceeds by normalizing each row of $M$ and then multiplying $\widetilde{M}$ by its transpose. Row normalization touches each entry exactly once and therefore costs $\Theta(Kd)$ time. The Gram product $\widetilde{M}\widetilde{M}^{\top}$ consists of $K^{2}$ dot products of length $d$, which is $\Theta(K^{2}d)$ time ([86]). The cosine matrix itself occupies $K^{2}$ entries. If it is retained after thresholding, it uses $\Theta(K^{2})$ space. If dropped right after graph construction, that $\Theta(K^{2})$ storage is only temporary.
@@
$T_{\text{refresh}}\;=\;\Theta(Kd)\;+\;\Theta(K^{2}d)\;+\;\Theta(K^{2})\;+\;O(K^{2})\;=\;\Theta(K^{2}d),$ (G.4)
@@
$S_{\text{refresh}}\;=\;\Theta(Kd)\;+\;\Theta(K^{2}),$ (G.5)
@@
G.3 Amortized cost over training
@@
$\frac{1}{T}\,N_{\text{refresh}}\,T_{\text{refresh}}\;\leq\;\frac{1}{T}\,\Big(\frac{T}{R}+1\Big)\,T_{\text{refresh}}\;=\;\frac{1}{R}\,T_{\text{refresh}}+\frac{1}{T}\,T_{\text{refresh}}$ (G.6)
@@
$\frac{1}{R}\,T_{\text{refresh}}\;=\;\ \frac{1}{R}\,\Theta(K^{2}d)\$(G.7)
@@
G.4 Conditions for negligible overhead
@@
$C_{\text{sched}}=\frac{a}{R}\,K^{2}d\qquad\text{and}\qquad C_{\text{probe}}=\frac{b}{R}\,K\,\mathsf{G},$
@@
$\frac{C_{\text{sched}}}{C_{\text{probe}}}=\frac{a}{b}\,\frac{K^{2}d}{K\,\mathsf{G}}=\frac{a}{b}\,\frac{Kd}{\mathsf{G}}.$
@@
$\frac{C_{\text{sched}}}{C_{\text{probe}}}\leq\varepsilon\quad\text{for some }0<\varepsilon\ll 1\quad\Longleftrightarrow\quad Kd\;\leq\;\frac{b}{a}\,\varepsilon\,\mathsf{G}.$
G.5 Reducing time complexity
@@
G.5.1 Random projections
We replace the EMA matrix $M\in\mathbb{R}^{K\times d}$ by a lower-dimensional sketch $\widetilde{M}=MR$ with $R\in\mathbb{R}^{d\times r}$ and $r\ll d$ ([87]). The sketching multiply costs $O(Kdr)$ and the cosine Gram becomes $O(K^{2}r)$ instead of $\Theta(K^{2}d)$. Storage for the sketched EMAs is $O(Kr)$. By the Johnson-Lendnstrauss (JL) random projection guarantee, if we map the $K$ task-EMA vectors from $\mathbb{R}^{d}$ to $\mathbb{R}^{r}$ using a suitable random matrix with $r=\Theta(\epsilon^{-2}\log K)$, then after row normalization all pairwise inner products (hence cosines) are preserved within $\pm\epsilon$ with high probability. We assume a uniform row-norm floor $\min_{i}\|m_{i}\|\geq m_{0}>0$ (which can be enforced in practice by skipping tasks with $\|m_{i}\|<\nu\ll m_{0}$) so cosine errors remain controlled. Choosing $\epsilon<\gamma$, where $\gamma$ is the cosine margin from the recovery analysis, ensures that every pair remains on the same side of the threshold $\tau$. Therefore the set $\{(i,j):\widehat{C}_{ij}<-\tau\}$ and the resulting coloring are unchanged with high probability.
@@ G.5.1 Random projections
G.5.2 Deterministic covariance sketching via frequent directions
We maintain a deterministic sketch $B\in\mathbb{R}^{\ell\times d}$ of the row space of $M$ using Frequent Directions and either project rows onto $\mathrm{span}(B)$ or form an approximate Gram from the sketch ([88, 89]). Maintaining the sketch costs $O(Kd\ell)$, the cosine Gram in the sketch space costs $O(K^{2}\ell)$, and storage for the sketch is $O(\ell d)$. Frequent Directions gives a spectral-norm bound
$|MM^{\top}-\widehat{MM^{\top}}|_{2}\leq\epsilon|M|_{F}^{2}$ (G.8)
@@ G.5.2 Deterministic covariance sketching via frequent directions
$\big|\cos(m_{i},m_{j})-\widehat{\cos}(m_{i},m_{j})\big|\;\leq\;\frac{2\,\epsilon\,\|M\|_{F}^{2}}{m_{0}^{2}}\;+\;O\!\left(\frac{\epsilon^{2}\,\|M\|_{F}^{4}}{m_{0}^{4}}\right)\$(G.9)
@@
G.5.3 Edge sampling for conflict graphs with adaptive refinement
We reduce the number of cosine evaluations by computing $\widehat{C}_{ij}$ for only $\tilde{O}(K\log K)$ randomly chosen task pairs to build a provisional conflict graph and then refining by evaluating additional pairs that are near the threshold or needed to certify connectivity and chromatic structure. We still compute all $K$ row norms once in $O(Kd)$ time for normalization, and the first pass costs $O(Kd\log K)$ for the sampled dot products. The total cost adds only the refinement work, which remains small when only few pairs are ambiguous. Under a planted separation model with margin $\gamma$ and reasonably dense cross-group conflicts, one can show with high probability that the sampled graph already captures the correct inter-group connectivity, so the coloring or component structure is recovered after the first pass and only boundary pairs need refinement. This reduces the pairwise work from $K^{2}$ to near $K\log K$ while preserving the final decisions under stated assumptions ([90]).
G.5.4 Incremental gram updates
@@
Appendix H Descent Bounds for Scheduled versus Aggregated Updates
$F(y)\ \leq\ F(x)+\langle\nabla F(x),y-x\rangle+\tfrac{L}{2}|y-x|^{2},\quad\forall x,y.$ (H.1)
H.1 Single refresh baselines and notation
H.1.1 Single aggregated step
Definition H.1 (Aggregated step).
$x^{\mathrm{agg}}\ :=\ x-\eta\,\sum_{r=1}^{m}G_{r}^{0}.$ (H.2)
One-shot $L$-smoothness bound.
Applying $L$-smoothness with $y=x^{\mathrm{agg}}$ yields
$F(x^{\mathrm{agg}})\ \leq\ F(x)\;-\;\eta\Big\langle\nabla F(x),\sum_{r=1}^{m}G_{r}^{0}\Big\rangle\;+\;\tfrac{L\eta^{2}}{2}\Big\|\sum_{r=1}^{m}G_{r}^{0}\Big\|^{2}.$ (H.3)
H.1.2 Scheduled group sequence over one refresh
Definition H.2 (Scheduled refresh).
Starting from the same point $x$, define
$x_{0}:=x,\qquad x_{r}:=x_{r-1}-\eta\,G_{r}(x_{r-1})\quad(r=1,\ldots,m),\qquad x^{\mathrm{sch}}:=x_{m}.$ (H.4)
Order and notation.
H.2 Telescoping bound for scheduled updates
Lemma 3 (Smoothness Expansion for Two Scheduled Groups).
Let $m=2$ and $G_{r}^{0}:=G_{r}(x)$. For any $\eta\in(0,1/L]$,
$F(x^{\text{sch}})$ $\leq F(x)\;-\;\eta\,\langle\nabla F(x),\,G_{1}^{0}\rangle\;+\;\tfrac{L\eta^{2}}{2}\,\|G_{1}^{0}\|^{2}$ (H.5)
Proof sketch.
$\nabla F(x_{1})=\nabla F(x)-\int_{0}^{1}H(x{-}t\eta G_{1}^{0})\,\eta G_{1}^{0}\,dt$ (H.6)
by the fundamental theorem of calculus along the segment $x\to x_{1}$. ∎
H.2.1 Start-of-refresh reduction under per-group lipschitzness
We adopt the following assumption whenever we compare bounds solely in terms of start-of-refresh measurements. It will be used throughout Sections H.3H.6
Assumption 6 (Per-group lipschitzness).
Each group map $G_{r}(\cdot)$ is $L_{r}$-lipschitz:
$\|G_{r}(u)-G_{r}(v)\|\ \leq\ L_{r}\,\|u-v\|\quad\text{for all }u,v.$ (H.7)
$\|G_{2}(x_{1})\|\leq\|G_{2}^{0}\|+L_{2}\eta\|G_{1}^{0}\|$ (H.8)
For general $m$
$\|G_{r}(x_{r-1})\|\ \leq\ \|G_{r}^{0}\|+L_{r}\,\eta\sum_{p<r}\|G_{p}^{0}\|\qquad(r=2,\dots,m)$ (H.9)
H.3 Upper bounds for scheduled and aggregated updates (general $m$)
Applying $L$-smoothness $m$ times yields the scheduled upper bound
@@ H.3 Upper bounds for scheduled and aggregated updates (general $m$)
$\mathrm{UB}_{\text{sch}}(x;\eta)$ $=F(x)\;-\;\eta\,\sum_{r=1}^{m}\big\langle\nabla F(x),\,G_{r}(x_{r-1})\big\rangle\;+\;\frac{L\eta^{2}}{2}\,\sum_{r=1}^{m}\|G_{r}(x_{r-1})\|^{2}$ (H.10)
The aggregated upper bound is the one-shot bound from Equation H.3, restated as

### v2

@@ Appendix D Computational Complexity of One Refresh (and Amortized Over Training)
We analyze the computational and memory complexity of the proposed interference-aware scheduler per refresh and its amortized cost over training. The former accounts for the cost of a single refresh operation while the latter represents the average cost distributed across all training steps. We distinguish the work required by the underlying multi-task training objective (e.g., backpropagation to obtain gradients) from the scheduler overhead (EMA maintenance, cosine computation, conflict graph construction, and color).
D.1 Notation
@@
D.2 Per-refresh complexity (time and space)
@@
$m_{i}\leftarrow\beta m_{i}+(1-\beta)g_{i}$ (65)
@@
$\widehat{C}=\widetilde{M}\widetilde{M}^{\top}$ (66)
where $\widetilde{M}$ is the row-normalized version of $M$. It thresholds $\widehat{C}$ at $\tau$ to obtain the conflict adjacency. Finally, it applies a graph-coloring routine to the resulting simple graph (Welsh & Powell, 1967).
@@
The construction of $\widehat{C}$ proceeds by normalizing each row of $M$ and then multiplying $\widetilde{M}$ by its transpose. Row normalization touches each entry exactly once and therefore costs $\Theta(Kd)$ time. The Gram product $\widetilde{M}\widetilde{M}^{\top}$ consists of $K^{2}$ dot products of length $d$, which is $\Theta(K^{2}d)$ time (Kågström et al., 1998). The cosine matrix itself occupies $K^{2}$ entries. If it is retained after thresholding, it uses $\Theta(K^{2})$ space. If dropped right after graph construction, that $\Theta(K^{2})$ storage is only temporary.
@@
$T_{\text{refresh}}\;=\;\Theta(Kd)\;+\;\Theta(K^{2}d)\;+\;\Theta(K^{2})\;+\;O(K^{2})\;=\;\Theta(K^{2}d),$ (67)
@@
$S_{\text{refresh}}\;=\;\Theta(Kd)\;+\;\Theta(K^{2}),$ (68)
@@
D.3 Amortized cost over training
@@
$\frac{1}{T}\,N_{\text{refresh}}\,T_{\text{refresh}}\;\leq\;\frac{1}{T}\,\Big(\frac{T}{R}+1\Big)\,T_{\text{refresh}}\;=\;\frac{1}{R}\,T_{\text{refresh}}+\frac{1}{T}\,T_{\text{refresh}}$ (69)
@@
$\frac{1}{R}\,T_{\text{refresh}}\;=\;\ \frac{1}{R}\,\Theta(K^{2}d)\$(70)
@@
D.4 Conditions for negligible overhead
@@
$C_{\text{sched}}=\frac{a}{R}\,K^{2}d\qquad\text{and}\qquad C_{\text{probe}}=\frac{b}{R}\,K\,\mathsf{G},$ (71)
@@
$\frac{C_{\text{sched}}}{C_{\text{probe}}}=\frac{a}{b}\,\frac{K^{2}d}{K\,\mathsf{G}}=\frac{a}{b}\,\frac{Kd}{\mathsf{G}}.$ (72)
@@
$\frac{C_{\text{sched}}}{C_{\text{probe}}}\leq\varepsilon\quad\text{for some }0<\varepsilon\ll 1\quad\Longleftrightarrow\quad Kd\;\leq\;\frac{b}{a}\,\varepsilon\,\mathsf{G}.$ (73)
D.5 Reducing time complexity
@@
D.5.1 Random projections
We replace the EMA matrix $M\in\mathbb{R}^{K\times d}$ by a lower-dimensional sketch $\widetilde{M}=MR$ with $R\in\mathbb{R}^{d\times r}$ and $r\ll d$ (Dasgupta & Gupta, 2003). The sketching multiply costs $O(Kdr)$ and the cosine Gram becomes $O(K^{2}r)$ instead of $\Theta(K^{2}d)$. Storage for the sketched EMAs is $O(Kr)$. By the Johnson-Lindenstrauss (JL) random projection guarantee, if we map the $K$ task-EMA vectors from $\mathbb{R}^{d}$ to $\mathbb{R}^{r}$ using a suitable random matrix with $r=\Theta(\epsilon^{-2}\log K)$, then after row normalization all pairwise inner products (hence cosines) are preserved within $\pm\epsilon$ with high probability. We assume a uniform row-norm floor $\min_{i}\|m_{i}\|\geq m_{0}>0$ (which can be enforced in practice by skipping tasks with $\|m_{i}\|<\nu\ll m_{0}$) so cosine errors remain controlled. Choosing $\epsilon<\gamma$, where $\gamma$ is the cosine margin from the recovery analysis, ensures that every pair remains on the same side of the threshold $\tau$. Therefore the set $\{(i,j):\widehat{C}_{ij}<-\tau\}$ and the resulting coloring are unchanged with high probability.
@@ D.5.1 Random projections
D.5.2 Deterministic covariance sketching via frequent directions
We maintain a deterministic sketch $B\in\mathbb{R}^{\ell\times d}$ of the row space of $M$ using Frequent Directions and either project rows onto $\mathrm{span}(B)$ or form an approximate Gram from the sketch (Liberty, 2013; Ghashami et al., 2016). Maintaining the sketch costs $O(Kd\ell)$, the cosine Gram in the sketch space costs $O(K^{2}\ell)$, and storage for the sketch is $O(\ell d)$. Frequent Directions gives a spectral-norm bound
$|MM^{\top}-\widehat{MM^{\top}}|_{2}\leq\epsilon|M|_{F}^{2}$ (74)
@@ D.5.2 Deterministic covariance sketching via frequent directions
$\big|\cos(m_{i},m_{j})-\widehat{\cos}(m_{i},m_{j})\big|\;\leq\;\frac{2\,\epsilon\,\|M\|_{F}^{2}}{m_{0}^{2}}\;+\;O\!\left(\frac{\epsilon^{2}\,\|M\|_{F}^{4}}{m_{0}^{4}}\right)\$(75)
@@
D.5.3 Edge sampling for conflict graphs with adaptive refinement
We reduce the number of cosine evaluations by computing $\widehat{C}_{ij}$ for only $\tilde{O}(K\log K)$ randomly chosen task pairs to build a provisional conflict graph and then refining by evaluating additional pairs that are near the threshold or needed to certify connectivity and chromatic structure. We still compute all $K$ row norms once in $O(Kd)$ time for normalization, and the first pass costs $O(Kd\log K)$ for the sampled dot products. The total cost adds only the refinement work, which remains small when only few pairs are ambiguous. Under a planted separation model with margin $\gamma$ and reasonably dense cross-group conflicts, one can show with high probability that the sampled graph already captures the correct inter-group connectivity, so the coloring or component structure is recovered after the first pass and only boundary pairs need refinement. This reduces the pairwise work from $K^{2}$ to near $K\log K$ while preserving the final decisions under stated assumptions (Erdős & Rényi, 1960).
D.5.4 Incremental gram updates
@@
Appendix E Descent Preservation Under $\tau$-Compatibility
E.1 Proof of Proposition 6
Proposition 6.
Let $S\subseteq\{1,\dots,K\}$ be a $\tau$-compatible task set. That is, every pair of gradients satisfies
$\bigl\langle g_{i},\,g_{j}\bigr\rangle\;\geq\;-\tau\,\lVert g_{i}\rVert\,\lVert g_{j}\rVert,\qquad\forall\,i\neq j\in S,0\leq\tau<1$ (76)
Then
$\Bigl\lVert\sum_{k\in S}g_{k}\Bigr\rVert^{2}\;\;\geq\;\;\bigl(1-\tau(|S|-1)\bigr)\sum_{k\in S}\lVert g_{k}\rVert^{2}.$ (77)
Proof.
We begin with the polarization identity for any finite set of vectors:
$\Bigl\lVert\sum_{k\in S}g_{k}\Bigr\rVert^{2}=\sum_{k\in S}\lVert g_{k}\rVert^{2}+2\!\!\sum_{\begin{subarray}{c}i,j\in S\\ i<j\end{subarray}}\!\bigl\langle g_{i},g_{j}\bigr\rangle.$ (78)
E.1.1 Lower-bounding the cross terms
Because $S$ is $\tau$-compatible, inequality (76) gives
$\bigl\langle g_{i},g_{j}\bigr\rangle\;\geq\;-\tau\,\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (79)
Insert this bound into (78) to obtain
$\Bigl\lVert\textstyle\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}\;-2\tau\!\!\sum_{i<j}\!\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (80)
E.1.2 Symmetrizing the mixed sum
Observe that
$\sum_{i<j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;=\;\frac{1}{2}\sum_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (81)
Substituting (81) into (80) yields
$\Bigl\lVert\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}\;-\;\tau\sum_{\begin{subarray}{c}i,j\\ i\neq j\end{subarray}}\lVert g_{i}\rVert\,\lVert g_{j}\rVert.$ (82)
E.1.3 Bounding the mixed sum via Cauchy-Schwarz
Apply the Cauchy-Schwarz inequality in $\mathbb{R}^{|S|}$ to the vectors $a=(\lVert g_{1}\rVert,\dots,\lVert g_{|S|}\rVert)$ and $\mathbf{1}=(1,\dots,1)$:
$\sum_{k}\lVert g_{k}\rVert=\langle a,\mathbf{1}\rangle\;\leq\;\lVert a\rVert\,\lVert\mathbf{1}\rVert=\Bigl(\sum_{k}\lVert g_{k}\rVert^{2}\Bigr)^{1/2}\,\sqrt{|S|}.$ (83)
Using $\left(\sum_{k}a_{k}\right)^{2}\leq|S|\sum_{k}a_{k}^{2}$ and (84),
$\sum_{i\neq j}\|g_{i}\|\,\|g_{j}\|\;=\;\Bigl(\sum_{k}\|g_{k}\|\Bigr)^{2}\;-\;\sum_{k}\|g_{k}\|^{2},$ (84)
we obtain the standard estimate
$\sum_{i\neq j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;\leq\;(|S|-1)\sum_{k}\lVert g_{k}\rVert^{2}.$ (85)
Hence,
$\tau\sum_{i\neq j}\lVert g_{i}\rVert\,\lVert g_{j}\rVert\;\leq\;\tau\,\bigl(|S|-1\bigr)\sum_{k}\lVert g_{k}\rVert^{2}.$ (86)
E.1.4 Combining bounds
Insert (86) into (82):
$\Bigl\lVert\sum_{k}g_{k}\Bigr\rVert^{2}\;\geq\;\sum_{k}\lVert g_{k}\rVert^{2}-\tau\bigl(|S|-1\bigr)\sum_{k}\lVert g_{k}\rVert^{2}\;=\;\bigl(1-\tau(|S|-1)\bigr)\sum_{k}\lVert g_{k}\rVert^{2},$ (87)
which is (77). ∎
E.2 Interpretation and practical implications
Equation (77) guarantees that whenever we restrict an SGD step to a $\tau$-compatible group (i.e., a set of tasks whose gradients are not too conflicting) the resulting joint update preserves at least a $\bigl(1-\tau(|S|-1)\bigr)$ fraction of the summed squared step lengths.
Below, we provide a strictly stronger version that is assumption free.
Proposition 7 (Data-Dependent Lower Bound via the Aggregate Conflict Ratio).
Define the aggregate conflict ratio
@@ E.2 Interpretation and practical implications
$\tau_{\mathrm{eff}}(S)\;:=\;\frac{\displaystyle\sum_{i\neq j}\bigl(-\langle g_{i},g_{j}\rangle\bigr)_{+}}{\displaystyle\sum_{k}\|g_{k}\|^{2}},\qquad(x)_{+}:=\max\{x,0\}.$ (88)
Then, without additional assumptions,
$\Bigl\lVert\sum_{k\in S}g_{k}\Bigr\rVert^{2}\;\geq\;\Bigl(1-\tau_{\mathrm{eff}}(S)\Bigr)\sum_{k\in S}\lVert g_{k}\rVert^{2},$ (89)
and under $\tau$-compatibility we always have $\tau_{\mathrm{eff}}(S)\leq\tau(|S|-1)$, so (89) is never weaker than (77).
Our takeaways from this are as follows:
  1. (i)
    Descent direction safety. The aggregated step is guaranteed to be a descent direction whenever $\tau_{\mathrm{eff}}(S)<1$ (data-dependent) and, in particular, whenever $\tau(|S|-1)<1$ (worst-case).
  2. (ii)
    Convergence-rate constant. In analyses for smooth SGD, one may replace $\lVert g_{t}\rVert^{2}$ by the right-hand side of either (89) (which is tighter) or (77) (worst-case), leading respectively to constants involving $\tau_{\mathrm{eff}}(S_{t})$ or $\tau(|S_{t}|-1)$.
Appendix F Convergence rate with $\tau$ -dependent constant
Theorem 7 (Baseline $O(1/\sqrt{T})$ convergence of the full gradient).
Let $F(\theta)=\sum_{k=1}^{K}\mathcal{L}_{k}(\theta,\phi_{k})$ be LL–smooth in the shared parameters $\theta$. Assume the stochastic gradient $g_{t}$ obtained at step $t$ satisfies $\mathbb{E}[g_{t}\mid\theta_{t}]=\nabla F(\theta_{t})$ and $\mathbb{E}[\lVert g_{t}-\nabla F(\theta_{t})\rVert^{2}\mid\theta_{t}]\leq\sigma^{2}$. Let the step size be $\eta=\tfrac{c}{\sqrt{T}}$ with $0<c\leq\frac{1}{L}$, and suppose the scheduler selects a $\tau$-compatible task set $S_{t}$ at each step (this will be used below for a refinement). Then
$\min_{1\leq t\leq T}\;\mathbb{E}\!\left[\lVert\nabla F(\theta_{t})\rVert^{2}\right]\;\leq\;\frac{2\bigl(F_{0}-F^{\star}\bigr)}{c\sqrt{T}}\;+\;\frac{cL\sigma^{2}}{\sqrt{T}}.$ (90)
Proof.
Because $F$ is LL–smooth, for any $\eta\leq\tfrac{1}{L}$ the standard non-convex SGD inequality (Ghadimi & Lan 2013, Lemma 3.2) gives
$\mathbb{E}\!\bigl[F(\theta_{t+1})\bigr]\;\leq\;\mathbb{E}\!\bigl[F(\theta_{t})\bigr]-\frac{\eta}{2}\,\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]+\frac{\eta^{2}L\sigma^{2}}{2}.$ (91)
Summing equation 91 over $t=0,\dots,T-1$ and using $\mathbb{E}[F(\theta_{T})]\geq F^{\star}$ yields
$\frac{\eta}{2}\sum_{t=0}^{T-1}\mathbb{E}\!\bigl[\lVert\nabla F(\theta_{t})\rVert^{2}\bigr]\;\leq\;F_{0}-F^{\star}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (92)
Dividing by $T$, using $\min_{t}x_{t}\leq\tfrac{1}{T}\sum_{t}x_{t}$, and substituting $\eta=\tfrac{c}{\sqrt{T}}$ gives equation 90. ∎
Data-dependent $\tau$-refinement for the scheduled gradient energy.
For a finite set $S$, define the aggregate conflict ratio